Sidney flew in an airplane 5 times during her vacation. The flights averaged 2 hours 28 minutes long. The lengths of her first four flights were 2 hours 10 minutes, 3 hours 20 minutes, 1 hour 15 minutes, and 2 hours 40 minutes. What was the length of her last flight?
Answer by TimothyLamb(4379)

x = length of 5th flight in hours
y = sum of lengths of first 4 flights in hours
---
y = (2+10/60) + (3+20/60) + (1+15/60) + (2+40/60)
y = 9.4166666666
---
( y + x )/5 = (2+28/60)
( y + x )/5 = 2.4666666666
x = 2.4666666666*5 - y
x = 2.4666666666*5 - 9.4166666666
x = 2.916666666
---
answer:
the 5th flight lasted 2.917 hours
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---